OpenStudy (anonymous):

MEDAL AND FAN!!! Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (anonymous):

\[-\frac{ 1 }{ 7 } ,-\frac{ 7 }{ 9 },-\frac{ 4 }{ 7 },9,-\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

Still doing imaginary test

OpenStudy (anonymous):

i beg your pardon?

OpenStudy (anonymous):

Nothing! ignore it

OpenStudy (anonymous):

ok but i actually need some help so unless you are going to help,please dont comment

OpenStudy (anonymous):

i have 2 more tests after this,a math and a global studies

OpenStudy (yanasidlinskiy):

Test? Exam? Please review the CoC. Thank you:)

OpenStudy (anonymous):

this itself is a worksheet type thing,the test is next